A plane traveling with the wind can travel 2400 miles in four hours. It takes six hours for the return flight. What is the speed of the plane in still air? what is the speed of the wind?
Answer by checkley77(12569)

2400/4=600mph
2400/6=400
(600+400)/2=1000/2=500 mph. is the air speed of the plane in still air.
(600-500)=100 mph. is the wind speed.
Proof:
400+100=500
600-100=500